# COSMOS: Cross-Modality Self-Distillation for Vision Language Pre-training

### v1

@@ 4.3 Zero-Shot Semantic Segmentation
As shown in Tab. 3 top, COSMOS trained on only 30M samples outperforms the best OpenCLIP model [13] trained on 1B samples in all benchmarks. In the datasets with a background category, COSMOS (29.6%, 12.5%, and 15.2%, respectively) surpasses the performance of OpenCLIP (1B) model (22.5%, 11.9%, and 11.4%, respectively) by a large margin. Remarkably, COSMOS almost doubles the performance on the fine-grained Cityscapes [15] dataset and achieves 20.0% average mIoU over all segmentation benchmarks compared to 16.5% obtained by the OpenCLIP model (1B). While CLIP-based models often suffer from feature suppression, COSMOS alleviates it through cross-modality embeddings and local-to-global matching.
In addition, we apply SCLIP [71] to our method where the weight of the last transformer block is adapted to construct the Correlative Self-Attention (CSA) block. Wang et al. [71] claimed that the CSA block maximizes the localization of patch embeddings without extra training. Although the CSA mechanism of SCLIP is tailored for the original CLIP model, it also improves the segmentation results of COSMOS. As a result, our COSMOS, trained on 30M samples, achieves an average mIoU of 37.8%, which is comparable to the 38.2% achieved by SCLIP [71], which was trained on 400M samples. Notably, COSMOS outperforms SILC [53] in several benchmarks despite a significantly smaller training budget (30M vs 10B samples). For instance, our model (vs SILC) achieves 34.7% and 23.2% (vs 26.9% and 20.8%) in terms of mIoU on the Cityscapes and COCO-stuff datasets. These results suggest that COSMOS can effectively generate fine-grained multi-modal representations with a much smaller dataset size.
@@ 4.4 Visual Perception & Contextual Understanding
We evaluate COSMOS on the visual perception and contextual understanding tasks [69, 33, 32] to demonstrate the robustness of our representations. COSMOS and previous methods [59, 80, 83] trained on the Merged-30M dataset are compared in Tab. 4, alongside the OpenCLIP [13] models trained on larger datasets. These tasks are specifically designed to assess the perception capabilities of vision-language models, as both visual and text input must be understood in detail to solve them.
Increasing the data size generally improves the performance of the OpenCLIP models [13], as shown by the SugarCrepe results. Interestingly, COSMOS not only surpasses the previous baselines, but also outperforms the OpenCLIP models, demonstrating that self-distillation and cross-modality learning play crucial roles in the compositional understanding of multi-modal data.
It is also noteworthy that COSMOS (trained on 30M samples) achieves the same performance (25.9%) as OpenCLIP [13] (trained on 1B samples) on the challenging benchmark MMVP-VLM [69]. This result suggests that COSMOS utilizes the pre-training data more efficiently, being able to learn distinguishing features with less data. Additionally, on the SugarCrepe dataset [33], where the task is to choose the correct caption among two captions (e.g., in the wrong caption some words are replaced, swapped inside the sentence or added), COSMOS achieves an average accuracy of 86.6%. This result signifies that COSMOS is able to differentiate between subtle differences in the text and, to a great extent, it does not treat the text as a bag-of-words.
@@ 4.5 Ablation Study
Table 5: Ablation. We report top-1 accuracy on ImageNet [18], and R@1 score for I2T and T2I retrieval tasks on Flickr30K [78] and MSCOCO [45]. All models are trained on CC3M with 1k batch size and ViT-B/16 image encoder.
@@
Additionally, we reproduce previous self-supervised methods based on CLIP, such as SILC [53] and SLIP [52], using the same synthetic dataset on which COSMOS is trained. We then directly compare these methods with COSMOS in the supplementary Sec. E.1.
@@ 4.6 Qualitative Results
We present additional qualitative results of COSMOS in Fig. 3. We visualize the normalized attention weights of the cross-attention modules for both images and text. For clarity, we use only three colors for the captions, discretized by their original weight values (dark blue for low value, light green for intermediate value and yellow for high value). Image attention results are based on the input text, while text attention results are conditioned on the input image.
The qualitative results illustrate that various objects are highlighted in both the image and text, regardless of their location and size. In the top-left example, the sheep from the foreground and turbines from the background are accurately captured in both image and caption, despite occupying only a small part of the image. Additionally, our module can attend to specific attributes of objects, such as the wooden fence or the silver sports car in the bottom-right example. More qualitative results are in the supplementary Sec. A.
Figure 3: Visualization of the attention map in the cross-attention modules. Attention weights are normalized between 0 and 1.
5 Conclusion and Limitations
Our COSMOS, a model for VL pre-training, uses cross-modality representation for self-distillation. Specifically, we design a text-cropping strategy based on long caption datasets, which enable us to learn local-to-global text correspondences. Our cross-attention module for self-supervised learning further encourages our model to learn both visual and textual grounding, overcoming previous shortcomings of the contrastive vision-language pre-training. Experimental results on various downstream tasks demonstrate that COSMOS outperforms strong baselines trained on much larger datasets on image-text retrieval, semantic segmentation and compositional understanding.
COSMOS is mainly restricted by the memory budget as it requires double inference during training. This may hinder the training of our method on larger batch sizes or models. Adjusting the number of global and local crops can mitigate this issue. Additionally, our text-cropping strategy is based on long captions generated by MLLMs, which may introduce hallucinations in the synthetic captions.
@@ 5 Conclusion and Limitations
Figure 4: Visualization of Attention Map. For different set of captions, we visualize the attention weights of image and text cross-attention modules. Patch-wise (image) and token-wise (caption) attention weights are both normalized between 0 and 1.
In this supplementary material, detailed information about the training procedure of COSMOS is included. We first provide additional qualitative results in Sec. A. The dataset configurations are described in Sec. B, and the experimental setups are outlined in Sec. C. Additionally, we elaborate on the baseline configurations in Sec. D and present further experiments in Sec. E.
@@ A Qualitative Results
In Fig. 4, we additionally visualize the attention maps of the cross-attention modules with different captions to illustrate their focus areas. We first normalize the attention weights across all patches and map them back onto the original image. Similarly, tokens in the captions are colored based on the normalized attention weights. The qualitative results demonstrate that our cross-attention module enhances the learning of local representations in both images and text. For example, our model is capable of detecting relatively small objects in the image, such as sugar cubes (first example of the second row) or a broken smartphone (third example of the first row). This information is often overlooked in previous models due to feature suppression.
@@ B Dataset Configuration
As COSMOS is trained on a web dataset, we had to download the images based on the corresponding URLs, but some of the URLs were invalid. Consequently, there is a discrepancy between the number of samples used to train our models and the original dataset sizes. The specific number of samples are reported in Tab. 6. We lost approximately 8.6% (YFCC15M) to 19.4% (CC12M) of samples compared to the original datasets due to expired URLs. This should be considered when directly comparing our models with others.
@@
The SugarCrepe dataset consists of images with positive and negative captions, where the model is required to choose the positive one given an input image (50% is the random chance). Negative captions are slightly different from the positive captions in terms of attributes and objects, with the goal of confusing the model. Since SugarCrepe is based on the COCO [45] validation set, we were able to download all images without any data loss.
SVO [32] consists of 36,841 data pairs, where each caption is paired with two images (one positive and one negative). Each image has its own triplet (subject, verb, object), and the positive and negative images differ in one component of the triplets. The original task is to match the caption with one of the two images based on similarity scores given by VLMs (50% by random chance). Since naive CLIP [59] already achieves 80% accuracy, we made this task harder by constructing negative captions. Based on the triplet of positive and negative images, we replaced the subject, verb, or object in the positive caption with the negative one. Currently, the model not only has to match the positive image with the positive caption but also match the negative image with the negative caption (25% by random chance). As SVO is released with image URL links, we were able to collect 22,220 data pairs (60.3% of the original size).
@@ C Experiment Configuration
In this section, we detail the configuration of our experiments, including hyperparameters for training (Sec. C.1), the inference process for each downstream task (Sec. C.2), and the pseudocode of our training procedure (Sec. C.3).
@@ C.1 Hyperparameters
As mentioned in the main paper, the teacher model is updated at each iteration using the Exponential Moving Average (EMA) of the student model. To obtain an effective teacher for self-distillation, we need to determine the momentum parameter $\lambda$, which controls the update rule of the teacher parameter $\theta_{t}$ based on the student parameter $\theta_{s}$ (i.e., $\theta_{t}=\lambda\theta_{t}+(1-\lambda)\theta_{s}$). According to Caron et al. [7], a higher batch size requires a lower momentum parameter. Based on their configuration, we choose 0.999 for CC3M and 0.99 for other datasets by default. We did not explicitly perform a grid search for this parameter. Empirically, we found that fixing the momentum parameter obtains a better performance, unlike Caron et al. [7], which used a cosine scheduler for the momentum parameter that converges to 1.0 eventually.
@@ C.2 Inference
Zero-shot retrieval is based on the image and text embeddings extracted from the pre-trained models, following the evaluation scheme in [76]. For each image-text pair, the pre-trained model generates image and text embeddings, collecting all embeddings for the entire dataset. We then compute a full cosine similarity matrix between all image embeddings and all text embeddings. Captions with the highest similarity scores are selected for each image to calculate the R@1, R@5, and R@10 metrics for image-to-text retrieval. Similarly, for text-to-image retrieval, we choose images with the highest scores for each caption. As each image in the validation or test set is equipped with multiple captions, image-to-text retrieval scores are generally higher than text-to-image scores.
In zero-shot semantic segmentation, we exclude the background category for PASCAL VOC [22] and PASCAL Context [51], following Cha et al. [8] and Wang et al. [71]. To clarify, we denote the original datasets with a background class as VOC21 and Context60, while the variants without the background category are referred to as VOC20 and Context59 in the main table. At inference time, for a given set of classes in the datasets, we obtain the corresponding text embeddings by querying our text encoder with a standard prompt. We compute the cosine similarity between the image patch embeddings (image tokens) and the text features of each class name to generate a segmentation map in a zero-shot manner. We adhere to the evaluation protocol established by SCLIP [71], including specifications for the window size, stride, and other parameters. We believe that the raw segmentation output of a VLM accurately reflects its zero-shot performance; therefore, we do not fine-tune our model or apply any post-refinement techniques such as PAMR [2].
For evaluation on SugarCrepe [33], SVO [32], and MMVP-VLM [69], we primarily referred to their evaluation demo prompts. Image and text embeddings are extracted from the pre-trained models, and then image-text pairs with higher cosine similarity scores are chosen as the final decision.
@@ C.3 Pseudocode
For clarity, we present the pseudocode of the training procedure of COSMOS in Algorithm 1. We empirically find that incorporating local image crops on the CLIP loss diminishes zero-shot performance, whereas using local text crops enhances it. Consequently, we compute the CLIP loss between all text crops and global image crops. We infer that integrating diverse captions during training boosts zero-shot transferability, while focusing on local image crops leads to overfitting on local details.
Algorithm 1 COSMOS: PyTorch-like Pseudocode
@@
h_img = s_cls_l + cross-attn(q=s_cls_l, kv=s_txt_tok_g)
h_txt = s_eot_l + cross-attn(q=s_eot_l, kv=s_img_tok_g)
@@ D Baseline Configuration
Table 8: OpenCLIP [13] model names
@@ E.1 Comparison to Previous SSL Methods
In Tab. 9, we directly compare COSMOS with the previous methods depicted in Fig. 5. For a fair comparison, we re-implemented these methods using the OpenCLIP [13] code repository and trained them on the CC3M dataset with long captions provided by DreamLIP [83]. The projection layer parameters follow the SLIP configuration, while the optimal temperature and loss scale for SILC were carefully determined. After a grid search, we selected student and teacher temperatures of 0.1 and 0.02, respectively, and used loss scale hyperparameters of (1.9, 0.1) for CLIP loss and self-distillation loss (see Tab. 16). Without extensive hyperparameter tuning, COSMOS significantly outperforms previous methods obtaining an accuracy of 37.1% in zero-shot classification on ImageNet [18] and 53.1% R@1 score in image-to-text retrieval on MSCOCO [45]. As mentioned in the introduction, focusing solely on enhancing image representation leads to sub-optimal results in VLMs, as shown by SILC which reached an accuracy score of only 30.4% on ImageNet and an R@1 score of 48.6% on image-to-text retrieval on MSCOCO, highlighting the importance of self-distillation with cross-modality representations.
@@
Table 9: Comparison to methods described in Fig. 5. All models are trained on CC3M with long synthetic caption. We use 1k batch size and ViT-B/16 image encoder.
@@ E.2 Experiments with ViT-B/32 Architecture
In the zero-shot classification tasks (Tab. 11), COSMOS improves all the metrics compared to CLIP [59] and SigLIP [80] across most datasets, while achieving comparable results to DreamLIP [83] on YFCC15M and Merged-30M. DreamLIP [83] aligns visual patch embeddings with positive text embeddings, promoting the learning of global structure by image tokens, which enhances classification performance when the model focuses on global information. Although COSMOS optimizes local-to-global correspondence to enhance the representation of local information through self-distillation, its performance is similar to DreamLIP. Additionally, increasing the dataset size generally improves accuracy, suggesting that the total number of images is a crucial factor for achieving high performance. The results obtained with ViT-B/32 are consistent with those observed using ViT-B/16 in the main paper.

### v2

@@ 4.3 Zero-Shot Semantic Segmentation
As shown in Tab. 3 top, COSMOS trained on only 30M samples outperforms the best OpenCLIP model [13] trained on 1B samples in all benchmarks. In the datasets with a background category, COSMOS (29.6%, 12.5%, and 15.2%, respectively) surpasses the performance of OpenCLIP (1B) model (22.5%, 11.9%, and 11.4%, respectively) by a large margin. Remarkably, COSMOS almost doubles the performance on the fine-grained Cityscapes dataset [15] and achieves 20.0% average mIoU over all segmentation benchmarks compared to 16.5% obtained by the OpenCLIP model (1B). While CLIP-based models often suffer from feature suppression, COSMOS alleviates this issue through cross-modality embeddings and local-to-global matching.
In addition, we apply SCLIP [77] to our method where the weight of the last transformer block is adapted to construct the Correlative Self-Attention (CSA) block. Wang et al. [77] claimed that the CSA block maximizes the localization of patch embeddings without extra training. Although the CSA mechanism of SCLIP is tailored for the original CLIP model, it also improves the segmentation results of COSMOS. As a result, our COSMOS, trained on 30M samples, achieves an average mIoU of 37.8%, which is comparable to the 38.2% achieved by SCLIP [77], which was trained on 400M samples. Notably, COSMOS outperforms SILC [57] in several benchmarks despite a significantly smaller training budget (30M vs 10B samples). For instance, our model (vs SILC) achieves 34.7% and 23.2% (vs 26.9% and 20.8%) on the Cityscapes and COCO-stuff datasets. These results suggest that COSMOS can effectively generate fine-grained multi-modal representations with a much smaller dataset size. In Sec. E.8, we present additional segmentation results comparing COSMOS to baselines trained on Merged-30M.
@@ 4.4 Visual Perception & Contextual Understanding
We evaluate COSMOS on the visual perception and contextual understanding tasks [74, 33, 32] to demonstrate the robustness of our representations. COSMOS and previous methods [63, 88, 91] trained on Merged-30M datasets are compared in Tab. 4, alongside the OpenCLIP [13] models trained on larger datasets. These tasks are specifically designed to assess the perception capabilities of Vision-Language Models, as both visual and text input must be understood in detail to solve them.
Increasing the data size generally improves the performance of OpenCLIP models [13] as shown by the SugarCrepe results. Interestingly, COSMOS not only surpasses the previous baselines, but also outperforms the OpenCLIP models, demonstrating that self-distillation and cross-modality learning play crucial roles in the compositional understanding of multi-modal data. It is also noteworthy that COSMOS (trained on 30M samples) achieves the same performance (25.9%) as OpenCLIP [13] (trained on 1B samples) on the challenging benchmark MMVP-VLM [74]. This result suggests that COSMOS utilizes the pre-training data more efficiently, being able to learn distinguishing features with less data. Additionally, on the SugarCrepe dataset [33], where the task is to choose the correct caption among two captions (e.g., in the wrong caption some words are replaced, swapped inside the sentence or added), COSMOS achieves an average accuracy of 86.6%. This result signifies that COSMOS is able to differentiate between subtle differences in the text and, to a great extent, it does not treat the text as a bag-of-words. In Sec. E.6, we presents results in the MLLM setting by integrating our vision encoder in the LLaVA framework [49].
@@ 4.5 Ablation Study
Table 5: Component ablation. We show top-1 accuracy on ImageNet [18], and R@1 score for I2T and T2I retrieval tasks on Flickr30K [84] and MSCOCO [48]. All models are trained on CC3M with the batch size of 1,024 and ViT-B/16 image encoder.
@@
Additionally, we reproduce previous self-supervised methods based on CLIP, such as SILC [57] and SLIP [56], using the same synthetic dataset on which COSMOS is trained. We directly compare these methods with COSMOS reporting the results in Sec. E.1.
@@ 4.6 Qualitative Results
We present additional qualitative results of COSMOS in Fig. 3. We visualize the normalized attention weights of the cross-attention modules for both images and text. For clarity, we use only three colors for the captions, discretized by their original weight values (dark blue for low value, light green for intermediate value, and yellow for high value). Image attention results are based on the input text, while text attention results are conditioned on the input image.
The qualitative results illustrate that various objects are highlighted in both the image and the text, regardless of their location and size. In the top-left example, the sheep from the foreground and turbines from the background are accurately captured in both image and caption, despite occupying only a small part of the image. Additionally, our module can attend to specific attributes of objects, such as the wooden fence or the silver sports car in the bottom-right example. More qualitative results are in Sec. A.
Figure 3: Visualization of attention map in cross-attention modules. Attention weights are normalized between 0 and 1.
5 Conclusion
Our COSMOS, a model for vision-language pre-training, uses cross-modality representation for self-distillation. Specifically, we design a text-cropping strategy based on long caption datasets, which enable us to learn local-to-global text correspondences. Our cross-attention module for self-supervised learning further encourages our model to learn both visual and textual grounding, overcoming the previous shortcomings of the contrastive vision-language pre-training. Experimental results on various downstream tasks demonstrate that COSMOS outperforms strong baselines trained on much larger datasets on image-text retrieval, semantic segmentation and compositional understanding.
@@ 5 Conclusion
Figure 4: Visualization of Attention Map. For different set of captions, we visualize the attention weights of the image and text cross-attention modules. The patch-wise (image) and token-wise (caption) attention weights are both normalized between 0 and 1.
In this supplementary material, we include detailed information about the training procedure of COSMOS. We first provide additional qualitative results in Sec. A. The dataset configurations are described in Sec. B, and the experimental setups are outlined in Sec. C. Additionally, we elaborate on the baseline configurations in Sec. D and present further experiments in Sec. E.
@@ A Qualitative Results
In Fig. 4, we visualize the attention maps of the cross-attention modules with different captions to illustrate their focus areas. We first normalize the attention weights across all patches and map them back onto the original image. Similarly, tokens in the captions are colored based on the normalized attention weights. The qualitative results show that our cross-attention module enhances the learning of local representations in both modalities, namely image and text. For example, our model is capable of detecting relatively small objects in the image, such as sugar cubes (first example of the second row) or a broken smartphone (third example of the first row). This information is often overlooked by the previous models due to feature suppression.
@@ B Dataset Configuration
As COSMOS is trained on web datasets, we had to download the images based on the corresponding URLs, but some of the URLs were invalid. Consequently, there is a discrepancy between the number of samples used to train our models and the original dataset sizes. The specific number of samples are reported in Tab. 6. We lost approximately 8.6% (YFCC15M) to 19.4% (CC12M) of samples compared to the original datasets due to expired URLs. This detail should be considered when directly comparing our models with others.
@@
The SugarCrepe dataset [33] consists of images with positive and negative captions, where the model is required to choose the positive one given an input image (50% is the random chance). Negative captions are slightly different from the positive captions in terms of attributes and objects, with the goal of confusing the model. Since SugarCrepe is based on the COCO [48] validation set, we were able to download all images without any data loss.
SVO [32] consists of 36,841 data pairs, where each caption is paired with two images (one positive and one negative). Each image has its own triplet (subject, verb, object), and the positive and negative images differ in one component of the triplets. The original task is to match the caption with one of the two images based on the similarity scores given by VLMs (50% by random chance). Since naive CLIP [63] already achieves 80% accuracy, we increased the difficulty of this task by constructing negative captions. Based on the triplet of positive and negative images, we replaced the subject, verb, or object in the positive caption with the negative one. Currently, the model not only has to match the positive image with the positive caption but also match the negative image with the negative caption (25% by random chance). As SVO is released with image URL links, we were able to collect 22,220 data pairs (60.3% of the original size).
@@ C Experiment Configuration
In this section, we detail the configuration of our experiments, including hyperparameters for training (Sec. C.1), the inference process for each downstream task (Sec. C.2), and the pseudocode of our training objective (Sec. C.3).
@@ C.1 Hyperparameters
As mentioned in the main paper, the teacher model is updated at each iteration using the exponential moving average (EMA) of the student model. To obtain an effective teacher for self-distillation, we need to determine the momentum parameter $\lambda$, which controls the update rule of the teacher parameter $\theta_{t}$ based on the student parameter $\theta_{s}$ (i.e., $\theta_{t}=\lambda\theta_{t}+(1-\lambda)\theta_{s}$). According to Caron et al. [7], a higher batch size requires a lower momentum parameter. Based on their configuration, we choose 0.999 for CC3M and 0.99 for other datasets by default. We did not explicitly perform a grid search for this parameter. Empirically, we found that fixing the momentum parameter obtains a better performance, unlike Caron et al. [7], who used a cosine scheduler for the momentum parameter which eventually converges to 1.0.
@@ C.2 Inference
Zero-shot retrieval is based on the image and text embeddings extracted from the pre-trained models, following the evaluation scheme presented in [82]. For each image-text pair, the pre-trained model generates image and text embeddings, collecting all embeddings for the entire dataset. We then compute a full cosine similarity matrix between all image embeddings and all text embeddings. Captions with the highest similarity scores are selected for each image to calculate the R@1, R@5, and R@10 metrics for image-to-text retrieval. Similarly, for text-to-image retrieval, we choose images with the highest scores for each caption. As each image in the validation or test set is equipped with multiple captions, image-to-text retrieval scores are generally higher than text-to-image scores.
In zero-shot semantic segmentation, we exclude the background category for PASCAL VOC [22] and PASCAL Context [55], following Cha et al. [8] and Wang et al. [77]. To be specific, we denote the original datasets with a background class as VOC21 and Context60, while the variants without the background category are referred to as VOC20 and Context59 in the main table. At inference time, for a given set of classes in the datasets, we obtain the corresponding text embeddings by querying our text encoder with a standard prompt. We compute the cosine similarity between the image patch embeddings (image tokens) and the text features of each class name to generate a segmentation map in a zero-shot manner. We adhere to the evaluation protocol established by SCLIP [77], including specifications for the window size, stride, and other parameters. We believe that the raw segmentation output of a VLM accurately reflects its zero-shot performance; therefore, we do not fine-tune our model or apply any post-refinement techniques such as PAMR [2].
To evaluate on SugarCrepe [33], SVO [32], and MMVP-VLM [74], we primarily referred to their evaluation demo prompts. Image and text embeddings are extracted from the pre-trained models, and then image-text pairs with higher cosine similarity scores are chosen as the final decision.
@@ C.3 Pseudocode
To increase the clarity of our method, we present the pseudocode of the training objective in Algorithm 1. We empirically found that incorporating local image crops on the CLIP loss diminishes zero-shot performance, whereas using local text crops enhances it. Consequently, we compute the CLIP loss between all text crops and global image crops. We infer that integrating diverse captions during training allows the model to learn various objects shown in the images, while including local image crops during training likely leads to the misalignment of image-text pairs.
Algorithm 1 COSMOS: Pseudocode of our loss function
@@
s_cls = {s_cls_g, s_cls_l} # Combine student crops
s_eot = {s_eot_g, s_eot_l} # Combine student crops
h_img = s_cls + cross-attn(q=s_cls, kv=s_txt_tok_g)
h_txt = s_eot + cross-attn(q=s_eot, kv=s_img_tok_g)
@@ D Baseline Configuration
Table 8: OpenCLIP [13] model names.
@@ E.1 Comparison to Previous SSL Methods
In Tab. 9, we directly compare COSMOS with the previous methods depicted in Fig. 5. For a fair comparison, we re-implemented these methods based on the OpenCLIP [13] code repository and trained them on the CC3M dataset with long captions provided by DreamLIP [91]. The projection layer parameters follow the SLIP configuration, while the optimal temperature and loss scale for SILC were carefully determined. After a grid search, we selected student and teacher temperatures of 0.1 and 0.02, respectively, and used loss scale hyperparameters of (1.9, 0.1) for CLIP loss and self-distillation loss (see Tab. 20). Without extensive hyperparameter tuning, COSMOS significantly outperforms previous methods obtaining an accuracy of 37.1% in zero-shot classification on ImageNet [18], and 53.1% R@1 score in image-to-text retrieval on MSCOCO [48]. As mentioned in the introduction, focusing solely on enhancing image representation leads to sub-optimal results in VLMs, as shown by SILC which reached an accuracy score of only 30.4% on ImageNet and an R@1 score of 48.6% on image-to-text retrieval on MSCOCO, highlighting the importance of self-distillation with cross-modality representations.
@@
Table 9: Comparison to methods described in Fig. 5. All models are trained on CC3M with long synthetic caption. We use the batch size of 1,024 and ViT-B/16 image encoder.
@@ E.2 Experiments with ViT-B/32 Architecture
In the zero-shot classification tasks (Tab. 11), COSMOS improves all metrics compared to CLIP [63] and SigLIP [88] across most datasets, while achieving comparable results to DreamLIP [91] on YFCC15M and Merged-30M. DreamLIP [91] aligns visual patch embeddings with positive text embeddings, promoting the learning of global structure by image tokens, which enhances the performance of classification task where the model focuses on global information. Although COSMOS optimizes local-to-global correspondence to enhance the representation of local information through self-distillation, its performance is similar to DreamLIP. Additionally, increasing the dataset size generally improves accuracy, suggesting that the total number of images is a crucial factor for achieving high performance in classification. The results obtained with ViT-B/32 are consistent with those of ViT-B/16 from the main paper.